Using the information below determine what government spending should be to obtain full employment.

-Full employment output = $90 million
-Exports = $10 million
-The marginal propensity to import = 0.15
-autonomous imports = $10 million
-The tax rate = 0.3
-Investment spending = $21 million
-Autonomous consumption = $17 million
-The marginal propensity to consume = 0.5
1
Expert's answer
2013-08-15T08:49:34-0400
Using the information below determine what government spending should be to
obtain full employment.


-Full employment output (Y) = $90 million
-Exports (E) = $10 million
-The marginal propensity to import (z') = 0.15
-autonomous imports (Z0) = $10 million
-The tax rate (T) = 0.3
-Investment spending (I) = $21 million
-Autonomous consumption (C0) = $17 million
-The marginal propensity to consume (c') = 0.5
- Government spending (G)=?
Solution
Full employment output (Y) is the quantity of real production or real aggregate output produced by the macroeconomy when resources are at full employment. It is estimated as the sum of consumer expenses (С), gross investments (I), government spending (G) and net flow of income into the country (NE):
Y'=C+G+I+NE,
Y'=Y-T*Y;
Consumer expenditures are all the purchases made by private individuals:
C=C0+c'*Y'.
Total gross investment includes all investment by consumers and businesses in securities or assets.
Exports minus imports represents the net flow of income into the country:
NE=E-Z=E-(Z0+z'*Y').
Government spending includes local, state and national expenditures. It is estimated as:
G=(Y-T*Y)-(C0+c'*(Y-T*Y))-I-(E-(Z0+z'*(Y-T*Y)))=90-0.3*90-(17+0.5*(90-0.3*90))-21-(10-(10+0.15*(90-0.3*90)))= $2.95 million
Answer $2.95 million